Shift report 20010815 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Wed Aug 15 2001 - 16:12:04 EDT

  • Next message: Apache: "Shift report 20010816 16:00-24:00"

    ------------------------------------------------------------
    Date: Wed Aug 15 16:12:03 2001, Shift: 08:00-16:00
    Supervisor: EunJoo, with: fv
    Log book#7 pages 91 - 94
    
    Shift Summary:
    ------------------------------------------------------------
    1. We had stable beam for a couple of hours.
       56 bunches.
    
    2. run 4210, 4211, 4214, 4216 - data run
       run 4212/4213 - pedestal run
    
    3. T3, T5 tripped while taking a data.
     
    4. brntuple is not running.
       (tried to update brop, but I don't know why...)
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Wed Aug 15 2001 - 16:13:20 EDT